Two rock climbers, Bill and Karen, use safety ropes of similar length. Karen's rope is more elastic, called a dynamic rope by climbers. Bill has a static rope, not recommended for safety purposes in pro climbing.

a)Karen falls freely about 2.1 m

m and then the rope stops her over a distance of 1.1 m

m . Estimate how large a force (assume constant) she will feel from the rope. (Express the result in multiples of her weight.)

b)In a similar fall, Bill's rope stretches by only 45 cm

cm . How many times his weight will the rope pull on him?

c)Which climber is more likely to be hurt?


1
Expert's answer
2021-11-22T10:13:37-0500

"a) \text{for \ Karen:}"

"P_k= m_kg-\text{weight Karen}"

"\text {the beginning of the fall:}"

"h_0= 0"

"v_0 = 0"

"E =EP = 0"

"\text{end of fall:}"

"h = -2.1m"

"\Delta x=-1.1m"

"h_1 = h+\Delta x=-3.2m"

"v_1=0,\text{hense }EK =0"

"E = EP_k+EP_r =0"

"EP_k-\text{Karen's potential energy}"

"EP_r\text{ -potential energy of the rope}"

"EP_k = m_kgh_1=-3.2m_kg"

"EP_r = -\frac{F_r \Delta x}{2}= 0.55F_r"

"-3.2m_kg+0.55F_r= 0"

"F_r = 5.82m_kg=5.82P_k"


"\text{Answer:}"

"\text{The force exerted by the rope is }5.82\text{ times the weight}"

"b)\text{ \text{for \ Bill:}}"

"\text{end of fall:}"

"h= -2.1m"

"\Delta x'=45cm =0.45m"

"h'_1 = h+\Delta x'=-2.55m"

"EP_b = m_bgh'_1=-2.55m_bg"

"EP'_r = -\frac{F'_r \Delta x'}{2}= 0.225F_r"

"-2.55m_bg+0.225F'_r= 0"

"F'_r = 11.33m_bg=11.33P_b"

"\text{Answer:}"

"\text{The force exerted by the rope is }11.33\text{ times the weight}"

"c)F_r =5.82P_k"

"F'_r =11.33P_b"

"a_k = 5.82g"

"a_b=11.33"

"\text{Bill's acceleration is greater than Karen's,}\newline
 \text{getting injured is also greater.}"

"\text{Answer:Bill is more likely to get hurt}"